A box contains 21 transistors, 7 of which are defective. If 7 are selected at random, find the probabpility that:
A. All are defective
B. None are defective

Answer by stanbon(57250)

A box contains 21 transistors, 7 of which are defective. If 7 are selected at random, find the probabpility that:
Binomial Problem with n = 21 ; P(defective) = 7/21 = 1/3 ; P(good) = 2/3
A. P(All are defective) = (1/3)^7 = 0.0004572...
------
B. P(None are defective) = (2/3)^7 = 0.0585
===================
Cheers,
Stan H.
================